# Decouple and Decompose: Scaling Resource Allocation with DeDe

### v2

@@ Abstract
Resource allocation is fundamental for cloud systems to ensure efficient resource sharing among tenants. However, the scale of such optimization problems has outgrown the capabilities of commercial solvers traditionally employed in production. To scale up resource allocation, prior approaches either tailor solutions to specific problems or rely on assumptions tied to particular workloads. In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: a vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand. Building on this insight, we develop DeDe, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a decouple-and-decompose approach. DeDe systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel. We have implemented DeDe as a library extension to an open-source solver, maintaining a familiar user interface. Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate DeDe’s substantial speedups and robust allocation quality.
@@ 1 Introduction
Resource allocation is critical to the orchestration of multi-tenant systems, which are managed by today’s cloud providers on an unprecedented scale. It allows cloud services to be delivered with efficiency, reliability, and cost-effectiveness, by judiciously distributing various compute resources (e.g., CPUs and GPUs) and I/O resources (e.g., memory, storage, and network bandwidth) across a vast number of clients.
Common applications of resource allocation include cluster scheduling [46, 27, 44, 61, 64], traffic engineering [24, 1, 67, 50, 31], and load balancing [16, 48, 52, 37]. To meet varying user or application demands (jobs, traffic flows, and workloads), these systems extensively employ commercial optimization solvers like Gurobi [22] to compute the allocation of expensive cloud infrastructure (cluster resources, network links, and servers) worth millions or even billions of dollars.
As cloud environments rapidly expand and diversify, the sheer scale of modern resource allocation problems has exceeded the capabilities of commercial solvers. These problems may involve millions of variables, which require solvers tens of minutes to hours to compute a solution when fast allocation is needed to maintain service quality. This “scalability crisis” in resource allocation has prompted a flurry of recent research [1, 43, 63, 47, 42] that trades off allocation quality for speed using heuristics, approximate algorithms, or machine learning. Nevertheless, these studies are limited to either a single domain (e.g., wide-area network traffic engineering [1, 63]) or a specific objective (e.g., max-min fairness [42]), requiring substantial manual effort to adapt to new scenarios.
In contrast, POP [43] takes a general approach to accelerating assorted resource allocation problems. It partitions a large problem into a few (e.g., 16) subproblems, each with a subset of resources and demands, and solves them independently. However, this partition hinges on a “granular” assumption—each demand requests only a small fraction of mutually interchangeable resources. This assumption proves brittle with real workloads and degrades solution quality (§7.2).
In this study, we re-examine real-world resource allocation problems through a different lens and aim to develop a general and scalable optimization framework without catering to specific problems or making workload-dependent assumptions. Our approach, DeDe, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
@@
This decomposition is made possible by an intrinsic, workload-agnostic problem structure that we uncovered after surveying dozens of real-world resource allocation problems in recent literature (Table 1). We discovered that a vast majority of these problems are inherently separable: they optimize the sum (or maximum) of utilities (measurement of solution quality) from allocating each individual resource and demand, while ensuring feasible allocation for individual resource and demand as well.111 We focus on individually separable resources and demands for simplicity and due to their widespread presence in real systems, but they can be relaxed to groups of resources and demands as we show in §4.1. In other words, the optimization objective comprises separate per-resource and/or per-demand utilities, and the constraints are also separate—they are either (1) a “resource constraint,” enforcing that the total demands mapped to each resource do not exceed the resource’s capacity, or (2) a “demand constraint,” enforcing that the total resources assigned to each demand do not surpass the demand’s needs. A formal definition of this separable structure is provided in §2, where we also show that real-world resource allocation is largely formulated with amenable mathematical properties (e.g., convex utilities, linear constraints) for tractability. [^2] Rare exceptions to this common structure are discussed in §4.2.
Nonetheless, the uncovered “separability” in resource allocation problems does not directly allow us to actually separate them into independent subproblems that address individual resources and demands as depicted in Figure 1. This is because each resource not only appears in its own resource constraint but also participates in the demand constraints associated with every demand mapped to it; likewise for each demand. Ignoring the dependencies across subproblems and greedily solving each subproblem independently would fail to find feasible and high-quality allocations.
Given the intertwined nature of resource and demand constraints, DeDe first decouples them as follows. Let $x$ denote the resource allocation matrix between $n$ resources and $m$ demands, where $x_{ij}$ represents the amount of resource $i$ assigned to demand $j$. The key mathematical trick in DeDe is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$. Next, DeDe iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed. In each iteration, DeDe further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$. This decoupling transformation lends itself to a classical constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9]. ADMM allows DeDe to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives. ADMM provides the theoretical foundation that enables alternating optimizations of these augmented Lagrangians with respect to $x$ and $z$, respectively.
Next, DeDe further decomposes the above optimizations into per-resource and per-demand allocations. This is feasible owing to the separable structure we identified across resource allocation problems: the original objective is separable on each resource and demand variable, and so are the penalty terms converted by ADMM from constraints in the augmented Lagrangians. After decomposition, per-resource allocations involve solving $n$ subproblems, each with $m$ demand variables, and $m$ subproblems with $n$ resource variables for per-demand allocations. Consequently, the scale of optimization is greatly reduced from $n\times m$ variables to only $n$ or $m$ variables in each subproblem. These subproblems can be solved independently with massive parallelism using off-the-shelf solvers. Figure 2 illustrates the “decouple and decompose” workflow of DeDe.
To demonstrate DeDe’s practicality, we have implemented it as a Python package, installable via pip install dede. DeDe is built on the widely used open-source solver cvxpy, and its interface is also designed to resemble cvxpy for ease of use (§6). Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized DeDe’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
@@
We evaluate DeDe in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7). Compared with the best variant of POP [43] for each problem, DeDe achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
Figure 2: Workflow of DeDe.
@@ 2 Real-World Resource Allocation Problems
In modern multi-tenant systems, resources such as CPUs, memory, and bandwidth are shared among numerous users and applications. A central resource allocator often casts the distribution of resources as an optimization problem (e.g., linear programming) and repeatedly computes feasible solutions with off-the-shelf solvers to accommodate changing demands.
However, these solvers are struggling to keep up with the increasing problem sizes that frequently occur in fast-growing systems. They may take several hours to solve an allocation problem that attempts to assign thousands of resources to thousands of demands—several orders of magnitude slower than desired (e.g., second-level SLAs, or service-level agreements). This challenge of scaling up resource allocation has spurred several recent works [1, 43, 63, 47, 42], but they either rely on domain-specific customizations or workload-dependent assumptions. In this paper, we instead seek to take a general (i.e., domain- and workload-agnostic) approach to scaling resource allocation.
After surveying real-world resource allocation problems from recent literature (Table 1), we discover that nearly all these formulations can be transformed into a separable structure (§1), which we formally describe as follows.
  • Variables.  An allocation matrix $x\in\mathcal{X}$ specifies how $n$ resources should be shared among $m$ demands, where $x_{ij}$ represents the quantity (or fraction) of resource $i$ assigned to demand $j$. Accordingly, the $i$th row of $x$, denoted as $x_{i*}=[x_{i1},x_{i2},\dots,x_{im}]$, constitutes the allocation vector of resource $i$, and the $j$th column $x_{*j}=[x_{1j},x_{2j},\dots,x_{nj}]$ encodes the allocation of demand $j$. The constraint set $\mathcal{X}$ is the Cartesian product of each $\mathcal{X}_{ij}$, which can be a set of real numbers or integers.
  • Objective.  The objective specifies a metric to be optimized over an allocation matrix $x$, primarily consisting of a sum of utilities (or costs, to measure allocation quality) from allocating each resource ($x_{i*}$) and demand ($x_{*j}$):
@@
    where $f_{i}(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}$ and $g_{j}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ are utility (or cost) functions. They are typically designed to be convex to ensure practical tractability. Minimizing the maximum of utilities can also be converted into regular minimizing utilities (Equation 1) with additional constraints [5].
  • Constraints.  To ensure that individual resources and demands are not over-provisioned (i.e., exceeding a resource’s capacity or a demand’s requirements), there is typically a constraint for each resource and each demand. We observe that these constraints are largely linear, presumably for tractability as well. For simplicity, we further convert inequality constraints into equality constraints using non-negative slack variables (details in Section 6). As such, the resource constraint on resource $i$ can be expressed as
    $R_{i}x_{i*}=r_{i},$ (2)
    and the demand constraint on demand $j$ can be written as
    $D_{j}x_{*j}=d_{j}.$ (3)
Despite the separable objective and separable constraints over each resource and each demand, large resource allocation problems (e.g., with millions of variables) still remain painfully slow to solve as they cannot actually be “separated” into smaller, concurrent per-resource and per-demand subproblems. The main obstacle is that resource and demand constraints are fundamentally coupled. Equations 2 and 3 show that each $x_{i,j}$ not only participates in the resource constraint (over $x_{i*}$) but also in the demand constraint (over $x_{*j}$). Next, we describe our solution to this challenge.
@@ 3 DeDe: Decouple and Decompose
We present DeDe, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems. Leveraging the separable structure (§2) that underpins a vast majority of these problems, DeDe enables massive parallelism with near-optimal solutions through a decouple-and-decompose technique. A formal, mathematical representation of DeDe is included in Appendix A.
@@ 3.1 Decouple
Resource and demand constraints being intertwined is the fundamental cause of complexity in resource allocation, as the individual allocation of any resource or demand will affect all other allocations. An optimal solution requires collectively allocating all resources to all demands, while a greedy strategy will lead to suboptimal or unfeasible solutions.
@@
Existing domain-agnostic solutions do not address this intrinsic problem nature of resource allocation. Production systems employ commercial solvers and rely on linear programming (LP) or mixed-integer linear programming (MILP) algorithms (e.g., simplex methods [45], barrier methods [30], or branch-and-bound methods [8]) to iteratively explore the feasible region—a high-dimensional search space dictated by coupled resource and demand constraints.
@@
The state of the art, POP [43], naively breaks intertwined constraints by partitioning resources and demands into a handful of subsets (e.g., 4–64), and then independently allocating each subset of resources to the demands within the same subset. Doing so ignores the dependencies across subsets and leads to suboptimal solutions in practice (§7.1). Meanwhile, it only enables limited parallelism because finer-grained decomposition would render the allocation unfeasible and unusable.
@@
Instead of disregarding the constraints across subproblems, we tackle the intertwining root cause directly. Specifically, we aim to decouple demand constraints from resource constraints, transforming the full allocation into alternating per-resource allocations—assigning a resource to all demands subject to the corresponding resource constraint, and per-demand allocations—mapping a demand to all resources are subject to the corresponding demand constraint.
To begin with, we offload demand constraints to another variable different from $x$. We create a copy of the allocation matrix $x$, denoted as $z$, and replace the variables in each demand constraint (Equation 3) with their counterparts in $z$:
$D_{j}z_{*j}=d_{j}.$ (4)
We also substitute $z$ for the demand-related utilities in the objective function (Equation 1):
$\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}).$ (5)
Then, representing the same allocation plan, $x$ and $z$ should be equal. Thus, we introduce a new constraint:
$x=z.$ (6)
Consequently, all resource constraints remain on $x$, but all demand constraints have been moved to $z$.
The decoupling of constraints clearly leads to a formulation equivalent to the original optimization with the new constraint $x=z$. However, the new constraint still has $x$ and $z$ entangled, so the introduction of $z$ does not immediately enable decomposition. In fact, jointly optimizing $x$ and $z$ (e.g., using penalty methods [6] or the augmented Lagrangian method [23]) will forfeit the benefits of decoupling, as we will show in §7.3.
On the bright side, the transformed formulation becomes compatible with a half-century-old constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9], which provides theoretically guaranteed solutions when solving for two sets (“blocks”) of variables ($x$ and $z$). At a high level, the theory behind ADMM enables alternating optimization of the following subproblems without making compromises on the solution quality: minimizing the updated objective (Equation 5) with respect to
  • $x$ (with $z$ fixed), subject to resource constraints and $x=z$,
  • $z$ (with $x$ fixed), subject to demand constraints and $x=z$.
We elaborate in §3.2 that both optimization subproblems are amenable to decomposition, reducing problem sizes and enabling massive parallelism. Next, we first describe how ADMM solves them (even without decomposition).
  • Constraint transformation.  ADMM is a variant of the augmented Lagrangian method [23]. Taking the first subproblem above (for $x$) as an example: ADMM converts every constraint, i.e., resource constraints and $x=z$, into penalty terms and appends them to the objective to minimize (Equation 5). A conventional penalty method will add a single penalty term $\sum_{i}\mu\cdot c_{i}^{2}$ to the objective, where $c_{i}$ is a measure of the constraint violation (comprising $|R_{i}x_{i*}-r_{i}|$ and $|x_{i}-z_{i}|$) for resource $i$, and $\mu$ is a penalty coefficient that increases to infinity over iterations to ensure asymptotic convergence to the optimal solution. However, taking $\mu$ to infinity requires careful manual tuning, or otherwise incurs ill-conditioning [2]. The augmented Lagrangian method has a similar rationale to penalty methods, except that it introduces yet another penalty term $\lambda_{i}\cdot c_{i}$ to mimic a Lagrange multiplier. The update of $\lambda_{i}$ is automated as $\lambda_{i}\leftarrow\lambda_{i}+\mu\cdot c_{i}$ with theoretical guarantees, where $\mu$ can stay as a small constant while avoiding ill-conditioning.
  • Partial optimization.  The new objectives with additional penalty terms converted from constraint violations are also called augmented Lagrangians. ADMM alternately minimizes these augmented Lagrangians, solving for $x$ with $z$ fixed, and solving for $z$ with $x$ fixed. This is known as “two-block partial optimization” and guarantees the convergence to optimal solutions for convex problems [21, 20]. For non-convex problems (including those with integer variables), practical evidence suggests that ADMM is still effective at solving (some of) them as well [53, 62, 68]. Theoretical proofs are available under specific assumptions [56, 38, 60].

### v3

@@ Abstract
Efficient resource allocation is essential in cloud systems to facilitate resource sharing among tenants. However, the growing scale of these optimization problems have outpaced commercial solvers commonly employed in production. To accelerate resource allocation, prior approaches either customize solutions for narrow domains or impose workload-specific assumptions. In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: the vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand. Building on this observation, we develop DeDe, a scalable and theoretically rooted optimization framework for large-scale resource allocation. At the core of DeDe is a decouple-and-decompose approach: it decouples entangled resource and demand constraints and thereby decomposes the overall optimization into alternating per-resource and per-demand subproblems that can be solved efficiently and in parallel. We have implemented and released DeDe as a Python package with a familiar modeling interface. Our experiments on three representative resource allocation tasks—cluster scheduling, traffic engineering, and load balancing—demonstrate that DeDe delivers significant speedups while generating higher-quality allocations.
This work is accepted at USENIX OSDI 2025.
@@ 1 Introduction
Resource allocation remains a critical and challenging problem today, especially as cloud providers operate multi-tenant systems on an unprecedented scale—these systems must ensure efficient and fair allocation of computing, storage, and networking resources across a large number of clients, in order to maintain the quality of cloud services.
Common scenarios of resource allocation include cluster scheduling [47, 28, 45, 63, 66], traffic engineering [24, 1, 69, 51, 32], and load balancing [14, 49, 53, 38]. To meet diverse and dynamic user demands (jobs, traffic flows, and queries), these systems extensively employ commercial optimization solvers like Gurobi [22] to solve linear programs (LP) or mixed-integer linear programs (MILP) that determine the allocation of valuable cloud resources (compute units, network links, and data nodes) worth millions to billions of dollars.
However, as cloud environments continue to expand and diversify, the sheer scale of modern resource allocation problems has exceeded the capabilities of commercial solvers. These problems may involve millions of variables and take tens of minutes or even hours to solve, whereas fast allocation is necessary to meet service-level objectives. This “scalability crisis” has spurred a flurry of studies [1, 65, 48, 43] that trade off solution quality for computational speed, using heuristics, approximate algorithms, or machine learning. Nevertheless, these approaches are typically restricted to specific domains (e.g., wide-area network traffic engineering) or particular optimization objectives (e.g., max-min fairness), limiting their broader applicability. While POP [44] accelerates assorted allocation tasks, it hinges on a “granular” assumption that each demand requests only a small fraction of interchangeable resources. This premise proves brittle under realistic workloads, leading to degraded solution quality (§7).
In this study, we revisit real-world resource allocation problems through a different lens, avoiding domain-specific designs or workload-dependent assumptions. We introduce DeDe  [^1] DeDe encapsulates our core technique “Decouple and Decompose”., a practical and scalable optimization framework for solving resource allocation problems. Rooted in optimization theory, DeDe allows the global allocation problem to be decomposed into alternating per-resource and per-demand subproblems (Figure 1), which can be solved more efficiently and in parallel without compromising solution quality.
@@
Figure 2: DeDe decouples constraints using an auxiliary variable, thereby decomposing optimization into subproblems.
This decomposition is enabled by an intrinsic, workload-agnostic problem structure that we identified after surveying dozens of real-world resource allocation problems in the literature (Table 1). We find that the vast majority of these problems are inherently separable: they optimize the total utility of individual resource and demand allocations, subject to separate constraints for each resource and demand to ensure feasibility. To be more concrete, we illustrate this problem structure using a simplified job scheduling example.
[…]
Formally, let $x_{ij}$ denote the fraction of the hour that job $j$ (demand) runs on GPU type $i$ (resource). Naturally, we have the demand constraints: $\forall j,\sum_{i}x_{ij}\leq 1$. On the other hand, for each GPU type, the total GPU hours requested by all jobs must not exceed the available GPU capacity. Thus, the resource constraints are: $\forall i,\sum_{j}\text{req}_{j}\cdot x_{ij}\leq\text{capacity}_{i}$. The objective is to maximize the total weighted average throughput, $\sum_{j}w_{j}\cdot(\sum_{i}\text{tput}_{ij}\cdot x_{ij})$, reflecting different job priorities. $\square$
This example demonstrates the “separable” structure frequently encountered in practical resource allocation problems: resource constraints are specified for each resource and its associated demands, demand constraints are specified for each demand and its contributing resources, and the objective function sums the utility achieved by each demand (or resource). Despite this seemingly modular structure, however, such optimization problems remain computationally challenging for commercial solvers to solve at scale. The difficulty arises from each allocation variable $x_{ij}$ simultaneously affecting both resource and demand constraints, precluding straightforward decomposition and parallel optimization.
To address this challenge, DeDe first decouples entangled resource and demand constraints through a mathematical reformulation. It introduces a copy of the allocation matrix $x$ as an auxiliary variable $z$, and rewrites the demand constraints (and per-demand terms in the objective) in terms of $z$, while adding a new constraint $x=z$. This transformation is mathematically equivalent, preserving the optimal solution. Moreover, it also lends itself to ADMM (Alternating Direction Method of Multipliers) [8], a constrained optimization method based on Lagrange multipliers [6]. By incorporating constraints into the objective via multipliers and constructing the (augmented) Lagrangian (see §3.1), ADMM provides a theoretical framework that allows DeDe to alternatively and iteratively optimize the Lagrangian with respect to each block of variables—$x$ or $z$—while holding the other fixed.
Building on the reformulation, DeDe further decomposes the optimization into subproblems along each resource and each demand. This step capitalizes on the identified separable structure, wherein the introduced Lagrangian can be partitioned across individual resources and demands (see §3.2). Consequently, instead of solving a monolithic optimization over all $n$ resources and $m$ demands, DeDe breaks the problem into $n$ per-resource subproblems, each with only $m$ demand variables, and $m$ per-demand subproblems, each with only $n$ resource variables. These smaller subproblems are amenable to independent and parallel optimization using off-the-shelf solvers. Figure 2 depicts the decouple-and-decompose workflow of DeDe.
To demonstrate DeDe’s practical utility, we have implemented DeDe as a Python package, installable via pip install dede. Built on the popular open-source modeling language cvxpy [15, 3], DeDe provides familiar APIs (§6) aligned with cvxpy. Moreover, unlike prior efforts that simulate parallelism [1, 44], our implementation is optimized for real parallel execution using Ray [41], enabling efficient utilization of many CPU cores with minimal overhead.
@@
We evaluate DeDe on three representative resource allocation tasks—cluster scheduling, traffic engineering, and load balancing, demonstrating its faster speed and higher allocation quality relative to the state of the art (§7). Compared with the best variant of POP [44] in each domain, DeDe achieves a 7.3% improvement in allocation quality and a 3.1$\times$ speedup for cluster scheduling, 5.3% and 7.6$\times$ for traffic engineering, and 12.6% and 2.2$\times$ for load balancing.
@@ 2 Real-World Resource Allocation Problems
In modern cloud systems, resources such as CPUs, memory, and network bandwidth are shared among numerous users and applications. A central resource allocator often casts the distribution of resources as an optimization problem (e.g., LP or MILP) and repeatedly computes feasible solutions using solvers to accommodate changing demands.
However, these solvers are struggling to keep up with the increasing problem sizes that frequently occur in fast-growing systems. They may take up to several hours to solve an allocation problem that attempts to assign thousands of resources to thousands of demands. This is several orders of magnitude slower than desired (e.g., service-level objectives measured in seconds). This challenge of scaling up resource allocation has prompted a series of recent works [1, 44, 65, 48, 43], but they either rely on domain-specific customizations or workload-dependent assumptions. In this work, we instead seek to take a general (i.e., domain- and workload-agnostic) approach to scaling resource allocation.
Separable problem structure.  Upon surveying real-world resource allocation problems from the literature (Table 1), we find that nearly all these formulations can be transformed into a separable structure (§1) formally described as follows. (We discuss the generality and limitations in §4.)
  • Variables.  An allocation matrix $x\in\mathcal{X}$ specifies the allocation of $n$ resources among $m$ demands, where $x_{ij}$ indicates the amount or fraction of resource $i$ assigned to demand $j$ (or vice versa). Accordingly, the $i$-th row of $x$, denoted as $x_{i*}=(x_{i1},x_{i2},\dots,x_{im})$, constitutes the allocation vector of resource $i$, while the $j$-th column $x_{*j}=(x_{1j},x_{2j},\dots,x_{nj})$ encodes the allocation of demand $j$. The constraint set $\mathcal{X}$ is the Cartesian product of the individual domains $\mathcal{X}_{ij}$, but it is often just non-negative real numbers or integers.
  • Objective.  The objective defines a metric to be maximized or minimized over an allocation matrix $x$, consisting of a sum of utilities or costs that quantify the outcome from allocating each resource $x_{i*}$ or demand $x_{*j}$:
@@
    where $f_{i}(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}$ and $g_{j}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ are utility or cost functions. These functions are typically convex (by nature or by choice), facilitating tractable optimization. A variety of objectives can be transformed into this form using standard optimization techniques. E.g., maximizing the minimum utility can be converted into maximizing an auxiliary “min utility” variable, with additional constraints ensuring that each utility is at least as large as this variable.
  • Constraints.  To enforce feasible allocation, one or more constraints can be imposed on each resource and each demand (e.g., to prevent oversubscription of resources and overprovisioning of demands). For computational tractability, these constraints are typically linear. Following standard optimization practices, we transform inequality constraints into equivalent equality constraints using slack variables. Thus, without loss of generality, we express the $rc_{i}\geq 1$ resource constraints on each resource $i$ as
    $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
    where $R_{i}\in\mathbb{R}^{rc_{i}\times m},r_{i}\in\mathbb{R}^{rc_{i}}$ are constraint parameters. Likewise, each demand $j$ is subject to $dc_{j}\geq 1$ linear constraints:
    $D_{j}x_{*j}=d_{j},\quad\forall j,$ (3)
    where $D_{j}\in\mathbb{R}^{dc_{j}\times n},d_{j}\in\mathbb{R}^{dc_{j}}$ are constraint parameters.
Despite the separable objective and separate constraints over each resource and each demand, large resource allocation problems (e.g., with millions of variables) still remain painfully slow to solve as they cannot actually be separated into smaller, concurrent subproblems. The main obstacle is that resource and demand constraints are fundamentally entangled. As shown in Equations 2 and 3, each $x_{i,j}$ not only participates in the resource constraints (over $x_{i*}$) but also in the demand constraints (over $x_{*j}$).
Figure 3: Toy scenario of the job scheduling example.
[…]
The left panel of Figure 3 shows the capacity of each GPU type ($\text{capacity}_{i}$) in GPU hours, alongside the throughput table ($\text{tput}_{ij}$). E.g., GPU type $C$ has a total capacity of 1.2 GPU hours, and assigning job 1 to it yields a throughput of 10 TPS.
The right panel shows the optimal allocation matrix ($x_{ij}$), which satisfies all resource and demand constraints. E.g., the optimal solution schedules job 1 on (one unit of) GPU type $A$ for 0.8 hours, and on GPU type $C$ for 0.2 hours. The maximum average throughput achieved is 18.8 TPS. $\square$
@@ 3 DeDe: Decouple and Decompose
We present DeDe, a practical and theoretically rooted optimization framework for solving large-scale, real-world resource allocation problems. By exploiting the inherently separable structure of these problems as described earlier, DeDe achieves massive parallelism while producing near-optimal solutions through a decouple-and-decompose approach.
@@ 3.1 Decouple
The interdependence between resource and demand constraints is a core complexity in resource allocation. Achieving an optimal solution requires jointly allocating all resources to all demands; otherwise, local, greedy strategies likely result in suboptimal or infeasible solutions.
@@
Existing solutions do not address this intertwined structure of constraints in resource allocation. In practice, production systems tend to simply rely on commercial solvers and employ linear programming (LP) or mixed-integer linear programming (MILP) algorithms (e.g., simplex methods [46], barrier methods [31], or branch-and-bound methods [7]) to navigate a high-dimensional search space, which is induced by the coupling of resource and demand constraints.
@@
A recent work, POP [44], naively breaks entangled constraints by partitioning resources and demands into a handful of subsets (e.g., 4–64), and then independently allocates a subset of resources to a subset of demands. Doing so ignores the dependencies across subsets and leads to suboptimal solutions in practice (§7). Meanwhile, it only enables limited parallelism because finer-grained decomposition would render the allocation infeasible and unusable.
@@
Instead of disregarding the constraints across subproblems, we tackle the root cause of constraint entanglement. Specifically, without sacrificing solution quality, our high-level goal is to decouple demand constraints from resource constraints, transforming the global allocation into alternating per-resource and per-demand subproblems.
Reformulation.  Consider the minimization of the objective in Equation 1 (the signs of $f_{i}$ and $g_{j}$ can be simply reversed for maximization problems), subject to resource constraints in Equation 2 and demand constraints in Equation 3.
First, we introduce an auxiliary variable $z$ as a duplicate of the allocation matrix $x$, and rewrite all demand constraints and all demand-related terms in the objective using $z$. Meanwhile, we add a new constraint $x=z$. This transformation clearly leads to an equivalent optimization problem:
$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}),$
$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (4)
$D_{j}z_{*j}=d_{j},\quad\forall j,$
$x-z=0.$
In this formulation, all resource constraints remain on $x$, while all demand constraints (and related utilities) are moved to $z$.
Although introducing $z$ appears to separate the problem structurally, the (necessary) constraint $x=z$ still tightly couples the variables and thus prevents straightforward decomposition. In fact, jointly optimizing $x$ and $z$ (e.g., using penalty methods [4] or augmented Lagrangian [23]) will forfeit the benefits of this reformulation, as we will show in §7.3.
ADMM background.  Nevertheless, this reformulation lends itself to ADMM (Alternating Direction Method of Multipliers) [8], a classical method for solving constrained optimization problems of the form: minimize $f(x)+g(z)$, subject to $Ax+Bz=c$. Note that Equation 3.1 conforms to this form.
ADMM is based on augmented Lagrangian [23], a method that converts constrained optimization into a sequence of unconstrained optimization problems using Lagrange multipliers [6] (a well-known technique in optimization). These unconstrained problems are defined over the Lagrangian function, constructed from the objective and constraints: $L_{\rho}(x,z,y)=f(x)+g(z)+y^{\mathsf{T}}(Ax+Bz-c)+(\rho/2)\lVert Ax+Bz-c\rVert^{2}_{2}$, where $y$ is (an estimate of) the Lagrange multiplier, and the last quadratic term with parameter $\rho$ serves as a penalty [4] for improving convexity and stability.
The theory of ADMM enables partial, alternating optimization of the Lagrangian with respect to $x$ and $z$—iteratively solving for $x$ with $z$ fixed, and solving for $z$ with $x$ fixed. This “two-block partial optimization” guarantees convergence to an optimal solution for convex problems [21, 20]. In non-convex settings (including those involving integers), empirical results indicate that ADMM can still effectively solve many of these problems [54, 64, 70]. Theoretical convergence results are established under certain conditions [57, 39, 61].
By introducing the auxiliary variable $u=(1/\rho)y$, the augmented Lagrangian can be further rewritten as: $L_{\rho}(x,z,u)=f(x)+g(z)+(\rho/2)\lVert Ax+Bz-c+u\rVert^{2}_{2}-(\rho/2)\lVert u\rVert^{2}_{2}$, leading to a more convenient scaled form of ADMM. Next, we present the ADMM iterates for updating $x$, $y$, and $u$ directly in the context of resource allocation problems.
Applying ADMM.  DeDe applies the scaled form of ADMM to solve the reformulated optimization problem (Equation 3.1). First, it constructs the augmented Lagrangian:
$\begin{split}&\mathcal{L}_{\rho}(x,z,\alpha,\beta,\lambda)=\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j})\\ &+\frac{\rho}{2}\Bigl{(}\sum_{i}\lVert R_{i}x_{i*}-r_{i}+\alpha_{i}\rVert_{2}^{2}+\sum_{j}\lVert D_{j}z_{*j}-d_{j}+\beta_{j}\rVert_{2}^{2}\\ &+\lVert x-z+\lambda\rVert_{F}^{2}\Bigr{)}-\frac{\rho}{2}(\lVert\alpha\rVert_{2}^{2}+\lVert\beta\rVert_{2}^{2}+\lVert\lambda\rVert_{F}^{2}),\end{split}$ (5)
where $\alpha_{i}\in\mathbb{R}^{rc_{i}}$, $\beta_{j}\in\mathbb{R}^{dc_{j}}$, $\lambda\in\mathbb{R}^{n\times m}$ are the introduced auxiliary variables. $\lVert\cdot\rVert_{2}$ still denotes the $\ell^{2}$-norm of a vector, while $\lVert\cdot\rVert_{F}$ is the Frobenius norm of a matrix, i.e., the $\ell^{2}$-norm of the matrix when flattened into a vector.
The ADMM iterates are given by:
$\alpha_{i}^{(k+1)}$ $:=\alpha_{i}^{(k)}+R_{i}x^{(k+1)}_{i*}-r_{i},\forall i$
$\beta_{j}^{(k+1)}$ $:=\beta_{j}^{(k)}+D_{j}z^{(k+1)}_{*j}-d_{j},\forall j$
$\lambda^{(k+1)}$ $:=\lambda^{(k)}+x^{(k+1)}-z^{(k+1)}$
We will show in §3.2 that both the optimization subproblems in Equations 6 and 7 are amenable to decomposition owing to the previously identified separable problem structure.